OpenStudy (lgbasallote):

Find the general solution: \[(r + \sin \theta - \cos \theta)dr + r(\sin \theta + \cos \theta)d\theta = 0\]

OpenStudy (lgbasallote):

is this exact?

OpenStudy (lgbasallote):

\[\frac{\partial M}{\partial \theta} = \cos \theta + \sin \theta\] \[\frac{\partial N}{\partial r} = \sin \theta + \cos \theta\] hmm it's exact...so i guess onto integration

OpenStudy (lgbasallote):

\[\frac{r^2}{2} + r\sin \theta - r\cos \theta + g(y)\]

OpenStudy (lgbasallote):

that's g(theta)

OpenStudy (lgbasallote):

\[r\cos \theta + r\sin \theta + g'(y) = r\sin \theta + r\cos \theta\] \[g(y) = 0\] so G.S> \[\frac{r^2}{2} + r\sin \theta - r\cos \theta= C\] correct?

OpenStudy (foolaroundmath):

Yeah it is correct.

OpenStudy (lgbasallote):

nice thanks